### v1

@@ Abstract
In this paper, we study second-order algorithms for the convex-concave minimax problem, which has attracted much attention in many fields such as machine learning in recent years. We propose a Lipschitz-free cubic regularization (LF-CR) algorithm for solving the convex-concave minimax optimization problem without knowing the Lipschitz constant. It can be shown that the iteration complexity of the LF-CR algorithm to obtain an $\epsilon$-optimal solution with respect to the restricted primal-dual gap is upper bounded by $\mathcal{O}(\frac{\rho\|z^{0}-z^{*}\|^{3}}{\epsilon})^{\frac{2}{3}}$, where $z^{0}=(x^{0},y^{0})$ is a pair of initial points, $z^{*}=(x^{*},y^{*})$ is a pair of optimal solutions, and $\rho$ is the Lipschitz constant. We further propose a fully parameter-free cubic regularization (FF-CR) algorithm that does not require any parameters of the problem, including the Lipschitz constant and the upper bound of the distance from the initial point to the optimal solution. We also prove that the iteration complexity of the FF-CR algorithm to obtain an $\epsilon$-optimal solution with respect to the gradient norm is upper bounded by $\mathcal{O}(\frac{\rho\|z^{0}-z^{*}\|^{2}}{\epsilon})^{\frac{2}{3}}$. Numerical experiments show the efficiency of both algorithms. To the best of our knowledge, the proposed FF-CR algorithm is the first completely parameter-free second-order algorithm for solving convex-concave minimax optimization problems, and its iteration complexity is consistent with the optimal iteration complexity lower bound of existing second-order algorithms with parameters for solving convex-concave minimax problems.
@@ 1 Introduction
where $f(x,y):\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable function. Throughout this paper, we assume that the function $f(x,y)$ is convex in $x$ for all $y\in\mathbb{R}^{n}$ and concave in $y$ for all $x\in\mathbb{R}^{m}$.
@@
Minimax optimization problems have a wide range of applications in game theory [23, 11], robust optimization [1, 38], and many other fields. Recently, new applications in the field of machine learning and data science, such as generative adversarial networks (GANs) [20, 30], adversarial learning [7], AUC maximization [21, 59], have further stimulated the research interest of many international scholars in the minimax problems.
@@
Depending on the available information of the objective function, there are three main types of optimization algorithms to solve the minimax problem (1), including zero-order, first-order, and second-order optimization algorithms, which use the function value, gradient, and Hessian information of the objective function, respectively. Compared with zero-order and first-order optimization algorithms, second-order optimization algorithms have attracted much attention due to their faster convergence speed. In this paper, we focus on second-order optimization algorithms for solving (1).
@@
Currently, there are relatively few studies on second-order algorithms for solving minimax optimization problems. Existing second-order algorithms can be divided into two categories, i.e., implicit algorithms [8, 34, 36, 39] and explicit algorithms [25, 26, 40, 47, 53].
@@
For the implicit case, these methods compute the subproblem of the current iteration involving update iteration, leading to implicit update rules. Monteiro and Svaiter [36] proposed a Newton proximal extragradient method for monotone variational inequality problems, which achieves an iteration complexity of $\mathcal{O}(\epsilon^{-2/3})$ in terms of the $\epsilon$-optimal solution of the primal-dual gap. To achieve faster convergence rates, methods that exploit $p$-order derivatives of $f(x,y)$ have been proposed, such as the higher-order generalized optimistic algorithm [39], the higher-order mirror-prox algorithm [8], and the restarted higher-order mirror-prox algorithms [34]. All of them can achieve an iteration complexity of $\mathcal{O}(\epsilon^{-2/(p+1)})$ in terms of the $\epsilon$-optimal solution of the primal-dual gap. In addition, [34] proved that the iteration complexity of the restarted higher-order mirror-prox algorithm is $\mathcal{O}(\epsilon^{-2/(p+1)}\log(\epsilon^{-1}))$ in terms of the $\epsilon$-optimal solution of the gradient norm.
@@
For the explicit case, Huang et al. [25] proposed a homotopy-continuation cubic regularized Newton algorithm for (1) with Lipschitz continuous gradients and Hessians, which achieves an iteration complexity of $\mathcal{O}(\log(\epsilon^{-1}))$ (resp. $\mathcal{O}(\epsilon^{-\frac{1-\theta}{\theta^{2}}})$ with $\theta\in(0,1)$) under a Lipschitz-type (resp. Hölderian-type) error bound condition in terms of the $\epsilon$-optimal solution of the gradient norm. For problems with only Lipschitz continuous Hessians, Lin et al. [47] proposed several Newton-type algorithms which all achieve an iteration complexity of $\mathcal{O}(\epsilon^{-2/3})$ in terms of the $\epsilon$-optimal solution of the restricted primal-dual gap. For more general monotone variational inequalities problems, Nesterov [53] proposed a dual Newton’s method with a complexity bound of $\mathcal{O}(\epsilon^{-1})$. Huang and Zhang [26] proposed an approximation-based regularized extra-gradient scheme, which exploits $p$-order derivatives of $f(x,y)$ and achieves an iteration complexity of $\mathcal{O}(\epsilon^{-2/(p+1)})$. All the above explicit methods need to solve cubic or higher-order regularized subproblems in each iteration.
@@
It should be noted that for most of the existing results mentioned above, achieving optimal complexity of the algorithm requires the assumption of knowing precise information about some parameters of the problem, such as the Lipschitz constant, the upper bound of the distance from the initial point to the optimal solution, etc. Accurately estimating these parameters is often challenging, and conservative estimates can significantly impact algorithm performance [16]. Therefore, designing algorithms with complexity guarantees without relying on inputs of these parameters has attracted considerable attention recently. This type of method is called parameter-free algorithm.
@@
There are very few existing research results on parameter-free second-order methods for solving minimax problems (1). Recently, Jiang et al. [40] proposed an adaptive second-order optimistic method, which solves a quadratic regularization subproblem at each iteration. It achieves an iteration complexity of $\mathcal{O}(\epsilon^{-2/3})$ (resp. $\mathcal{O}(\epsilon^{-1})$) in terms of the $\epsilon$-optimal solution of the restricted primal-dual gap (resp. gradient norm) under the assumption that the Hessian is Lipschitz continuous when the Lipschitz constant is known. If the Lipschitz constant is not known, the algorithm then requires an additional assumption that the gradient is Lipschitz continuous and can prove similar iteration complexity. Liu and Luo [10] proposed a $\nu$-regularized extra-Newton method for solving monotone variational inequality problems with an iteration complexity of $\mathcal{O}(\epsilon^{-\frac{2}{2+\nu}})$ when $\nu$ is known, and proposed an universal regularized extra-Newton method with an iteration complexity of $\mathcal{O}(\epsilon^{-\frac{4}{3(1+\nu)}})$ when $\nu$ is unknown, where $\nu$ is the Hölder continuous constant. Although [10] and [40] eliminate the need for prior information on the Lipschitz or Hölder parameters, they can only be considered partially parameter-free. This is because the algorithms require prior information such as a predetermined maximum number of iterations $T$, bounds on the constraint set, and the optimal solution, etc.
@@
To the best of our knowledge, there were no fully parameter-free optimization algorithms for solving minimax problems before. In this paper, we propose a fully parameter-free second-order algorithm for solving (1) that achieves optimal complexity without a prior knowledge of any parameters.
@@ 1.1 Other Related Works
There are many existing research results on first-order optimization algorithms for solving convex-concave minimax optimization problems. For instance, Korpelevich [18] proposed an extra-gradient (EG) algorithm, and its linear convergence rate for solving smooth strongly-convex-strongly-concave minimax problems and bilinear minimax problems was proved in [35]. Nemirovski [5] proposed a mirror-prox method which achieves the optimal iteration complexity, i.e., $\mathcal{O}(\epsilon^{-1})$, among algorithms for solving convex-concave minimax problems using only first-order information [58]. Nesterov [54] proposed a dual extrapolation algorithm, which has the same optimal iteration complexity. Mokhtari et al. [3, 4] established an overall complexity of $\mathcal{O}(\log\epsilon^{-1})$ (resp. $\mathcal{O}(\epsilon^{-1})$) for bilinear and strongly-convex-strongly-concave minimax problems (resp. smooth convex-concave minimax problems) for both the EG and the optimistic gradient descent ascent (OGDA) [28] method. For more related results, we refer to [57, 32, 17] and the references therein.
@@
For nonconvex-strongly-concave minimax problems, various first-order algorithms have been proposed in recent works [37, 14, 9, 44, 45, 42, 19], and all of them can achieve the iteration complexity of $\tilde{\mathcal{O}}\left(\kappa_{y}^{2}\epsilon^{-2}\right)$ in terms of stationary point of $\Phi(\cdot)=\max_{y\in\mathcal{Y}}f(\cdot,y)$ (when $\mathcal{X}=\mathbb{R}^{m}$), or stationary point of $f(x,y)$, where $\kappa_{y}$ is the condition number for $f(x,\cdot)$. Furthermore, Zhang et al. [43] proposed a generic acceleration framework which can improve the iteration complexity to $\mathcal{O}\left(\sqrt{\kappa_{y}}\epsilon^{-2}\right)$. Some second-order algorithms have also been proposed in recent works [29, 49, 60], all of them can achieve the iteration complexity of $\tilde{\mathcal{O}}(\epsilon^{-3/2})$ to find an $\epsilon$-second-order stationary point.
@@
For general nonconvex-concave minimax problems, there are two types of algorithms, i.e., multi-loop algorithms and single-loop algorithms. Various multi-loop algorithms have been proposed in [50, 45, 31, 13, 19, 27]. The best known iteration complexity of multi-loop algorithms for solving nonconvex-concave minimax problems is $\tilde{\mathcal{O}}(\epsilon^{-2.5})$, which was achieved by [45]. For solving nonconvex-linear minimax problems, [51] proposed a new alternating gradient projection algorithm with the iteration complexity of $\mathcal{O}(\epsilon^{-3})$. Various single-loop algorithms have also been proposed, e.g., the gradient descent-ascent (GDA) algorithm [44], the hybrid block successive approximation (HiBSA) algorithm [42], the unified single-loop alternating gradient projection (AGP) algorithm [64], and the smoothed GDA algorithm [24]. Both the AGP algorithm and the smoothed GDA algorithm achieve the best known iteration complexity, i.e., $\mathcal{O}(\epsilon^{-4})$, among single-loop algorithms for solving nonconvex-concave minimax problems.
@@
There are also some research results on the zero-order method for solving minimax problems. For the nonconvex-strongly-concave setting, Liu et al. [41] proposed a ZO-Min-Max algorithm with $\mathcal{O}(\kappa_{y}^{6}\epsilon^{-6})$ iteration complexity, where $\kappa_{y}$ is the condition number for $f(x,\cdot)$. Wang et al. [61] proposed a zeroth-order SGDA algorithm and a zeroth-order gradient descent multi-step ascent algorithm, and the total complexity to find an $\epsilon$-stationary point of $\Phi(\cdot)=\max_{y\in\mathcal{Y}}f(\cdot,y)$ is $\mathcal{O}\left(\kappa_{y}^{5}\epsilon^{-4}\right)$ and $\tilde{\mathcal{O}}\left(\kappa_{y}^{2}\epsilon^{-4}\right)$, respectively. Huang et al. [14] proposed an accelerated zeroth-order momentum descent ascent method with $\tilde{\mathcal{O}}(\kappa_{y}^{4.5}\epsilon^{-3})$ iteration complexity. Xu et al. [63] proposed a zero-order alternating stochastic gradient projection (ZO-AGP) algorithm to solve the nonconvex-concave minimax problem, and proved that the iteration complexity of the ZO-AGP algorithm is $\mathcal{O}(\epsilon^{-4})$. Xu et al. [62] proposed a zeroth-order alternating gradient descent ascent (ZO-AGDA) algorithm and a zeroth-order variance reduced alternating gradient descent ascent (ZO-VRAGDA) algorithm for solving a class of nonconvex-nonconcave minimax problems under the deterministic and the stochastic setting, respectively. The iteration complexity of the ZO-AGDA algorithm and the ZO-VRAGDA algorithm have been proved to be $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-3})$, respectively. For more related results, we refer to [2, 6, 48, 22] and the references therein.

### v2

@@ Abstract
In this paper, we study second-order algorithms for the convex-concave minimax problem, which has attracted much attention in many fields such as machine learning in recent years. We propose a Lipschitz-free cubic regularization (LF-CR) algorithm for solving the convex-concave minimax optimization problem without knowing the Lipschitz constant. It can be shown that the iteration complexity of the LF-CR algorithm to obtain an $\epsilon$-optimal solution with respect to the restricted primal-dual gap is upper bounded by $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{2}\epsilon^{-2/3})$ , where $z_{0}=(x_{0},y_{0})$ is a pair of initial points, $z^{*}=(x^{*},y^{*})$ is a pair of optimal solutions, and $\rho$ is the Lipschitz constant. We further propose a fully parameter-free cubic regularization (FF-CR) algorithm that does not require any parameters of the problem, including the Lipschitz constant and the upper bound of the distance from the initial point to the optimal solution. We also prove that the iteration complexity of the FF-CR algorithm to obtain an $\epsilon$-optimal solution with respect to the gradient norm is upper bounded by $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{4/3}\epsilon^{-2/3})$. Numerical experiments show the efficiency of both algorithms. To the best of our knowledge, the proposed FF-CR algorithm is a completely parameter-free second-order algorithm, and its iteration complexity is currently the best in terms of $\epsilon$ under the termination criterion of the gradient norm.
Keywords: parameter-free, second-order algorithm, convex-concave minimax optimization
@@ 1 Introduction
where $f:\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable function. Throughout this paper, we assume that the function $f(x,y)$ is convex in $x$ for all $y\in\mathbb{R}^{n}$ and concave in $y$ for all $x\in\mathbb{R}^{m}$.
@@
Minimax optimization problems have a wide range of applications in game theory (Neumann and Morgenstern, 1953; Blackwell and Girshick, 1979), robust optimization (Ben-Tal et al., 2009; Gao and Kleywegt, 2023), and many other fields. Recently, new applications in the field of machine learning and data science, such as generative adversarial networks (GANs) (Goodfellow et al., 2014; Arjovsky et al., 2017), adversarial learning (Sinha et al., 2018), AUC maximization (Hanley and McNeil, 1982; Ying et al., 2016), have further stimulated the research interest of many international scholars in the minimax problems.
@@
Depending on the available information of the objective function, there are three main types of optimization algorithms to solve the minimax problem (1), including zeroth-order, first-order, and second-order optimization algorithms, which use the function value, gradient, and Hessian information of the objective function, respectively. Compared with zeroth-order and first-order optimization algorithms, second-order optimization algorithms have attracted much attention due to their faster convergence speed. In this paper, we focus on second-order optimization algorithms for solving (1).
@@
Currently, there are relatively few studies on second-order algorithms for solving minimax optimization problems. Existing second-order algorithms can be divided into two categories, i.e., implicit algorithms (Bullins and Lai, 2022; Ostroukhov et al., 2020; Monteiro and Svaiter, 2012; Jiang and Mokhtari, 2022) and explicit algorithms (Huang et al., 2022; Huang and Zhang, 2025; Jiang et al., 2024; Lin et al., 2022; Nesterov, 2006).
@@
For the implicit case, these methods compute the subproblem of the current iteration involving update iteration, leading to implicit update rules. Monteiro and Svaiter (2012) proposed a Newton proximal extragradient method for monotone variational inequality problems, which achieves an iteration complexity of $\mathcal{O}(\epsilon^{-2/3})$. To achieve faster convergence rates, methods that exploit $p$-order derivatives of $f(x,y)$ have been proposed, such as the higher-order generalized optimistic algorithm (Jiang and Mokhtari, 2022), the higher-order mirror-prox algorithm (Bullins and Lai, 2022), and the restarted higher-order mirror-prox algorithms (Ostroukhov et al., 2020). All of them can achieve an iteration complexity of $\mathcal{O}(\epsilon^{-2/(p+1)})$ in terms of the $\epsilon$-optimal solution of the primal-dual gap. In addition, Ostroukhov et al. (2020) proved that the iteration complexity of the restarted higher-order mirror-prox algorithm is $\mathcal{O}(\epsilon^{-2/(p+1)}\log(\epsilon^{-1}))$ in terms of the $\epsilon$-optimal solution of the gradient norm.
@@
For the explicit case, Huang et al. (2022) proposed a homotopy-continuation cubic regularized Newton algorithm for (1) with Lipschitz continuous gradients and Hessians, which achieves an iteration complexity of $\mathcal{O}(\log(\epsilon^{-1}))$ (resp. $\mathcal{O}(\epsilon^{-(1-\theta)\theta^{2}})$ with $\theta\in(0,1)$) under a Lipschitz-type (resp. Hölderian-type) error bound condition in terms of the $\epsilon$-optimal solution of $\|z_{k}-z^{*}\|$. For problems with only Lipschitz continuous Hessians, Lin et al. (2022) proposed several Newton-type algorithms which all achieve an iteration complexity of $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{2}\epsilon^{-2/3})$ in terms of the $\epsilon$-optimal solution of the restricted primal-dual gap. For more general monotone variational inequalities problems, Nesterov (2006) proposed a dual Newton’s method with a complexity bound $\mathcal{O}(\epsilon^{-1})$. Huang and Zhang (2025) proposed an approximation-based regularized extra-gradient scheme, which exploits $p$-order derivatives of $f(x,y)$ and achieves an iteration complexity of $\mathcal{O}(\epsilon^{-2/(p+1)})$. All the above explicit methods need to solve cubic or higher-order regularized subproblems in each iteration.
@@
It should be noted that for most of the existing results mentioned above, achieving optimal complexity of the algorithm requires the assumption of knowing precise information about some parameters of the problem, such as the Lipschitz constant, the upper bound of the distance from the initial point to the optimal solution, etc. Accurately estimating these parameters is often challenging, and conservative estimates can significantly impact algorithm performance (Lan et al., 2023). Therefore, designing algorithms with complexity guarantees without relying on inputs of these parameters has attracted considerable attention recently. This type of method is called parameter-free algorithm.
@@
There are very few existing research results on parameter-free second-order methods for solving minimax problems (1). Jiang and Mokhtari (2022) proposed an optimistic second-order method (OSOM), where a line search strategy is employed to remove the need for prior knowledge of the Lipschitz constant. It achieves an iteration complexity of $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{2}\epsilon^{-2/3})$ in terms of the $\epsilon$-optimal solution of the restricted primal-dual gap. Liu and Luo (2022) proposed a $\nu$-regularized extra-Newton method (ReNewton) for solving monotone variational inequality problems with an iteration complexity of $\mathcal{O}(\epsilon^{-2/(2+\nu)})$ when $\nu$ is known, and proposed an universal regularized extra-Newton method with an iteration complexity of $\mathcal{O}(\epsilon^{-4/(3+3\nu)})$ when $\nu$ is unknown, where $\nu$ is the Hölder continuous constant. The proposed methods also employ a line search strategy to eliminate the need for problem parameters. Recently, Jiang et al. (2024) proposed an adaptive second-order optimistic method (ASOM), which solves a quadratic regularization subproblem at each iteration. It achieves an iteration complexity of $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{2}\epsilon^{-2/3})$ (resp. $\mathcal{O}(\rho\|z_{0}-z^{*}\|^{2}\epsilon^{-1})$) in terms of the $\epsilon$-optimal solution of the restricted primal-dual gap (resp. gradient norm) under the assumption that the Hessian is Lipschitz continuous when the Lipschitz constant is known. If the Lipschitz constant is not known, the algorithm then requires an additional assumption that the gradient is Lipschitz continuous and can prove an iteration complexity of $\mathcal{O}\left((\ell^{4/3}\rho^{2/3}\|z_{0}-z^{*}\|^{2/3}+\rho^{2}\|z_{0}-z^{*}\|^{2})\epsilon^{-2/3}\right)$ $\left(\text{resp.}\leavevmode\nobreak\ \mathcal{O}\left((\ell\rho\|z_{0}-z^{*}\|+\rho^{2}\|z_{0}-z^{*}\|^{2})\epsilon^{-1}\right)\right)$ in terms of the $\epsilon$-optimal solution of the restricted primal-dual gap (resp. gradient norm).
@@
To the best of our knowledge, no parameter-free second-order algorithm has previously achieved the complexity bound $\mathcal{O}(\epsilon^{-2/3})$ with respect to the gradient norm for solving unconstrained convex-concave minimax problems. In this paper, we propose a completely parameter-free second-order algorithm for solving (1) that achieves this complexity without any prior knowledge of the parameters.
@@ 1.1 Other Related Works
There are many existing research results on first-order optimization algorithms for solving convex-concave minimax optimization problems. For instance, Korpelevich (1976) proposed an extra-gradient (EG) algorithm, and its linear convergence rate for solving smooth strongly convex-strongly concave minimax problems and bilinear minimax problems was proved in Tseng (1995). Nemirovski (2004) proposed a mirror-prox method, which achieves an iteration complexity of $\mathcal{O}(\ell\epsilon^{-1})$ in terms of the primal-dual gap for constrained minimax problems, where $\ell$ is the gradient Lipschitz continuous constant of $f(\cdot,\cdot)$. Nesterov (2007) proposed a dual extrapolation algorithm, which has the same iteration complexity. Mokhtari et al. (2020a) (Mokhtari et al. (2020b)) established an overall complexity of $\mathcal{O}({\color[rgb]{0,0,1}\kappa}\log\epsilon^{-1})$ (resp. $\mathcal{O}(\ell\epsilon^{-1})$) for bilinear and strongly convex-strongly concave minimax problems (resp. smooth convex-concave minimax problems) for both the EG and the optimistic gradient descent ascent (OGDA) method (Popov, 1980) , where $\kappa$ is the condition number of $f(\cdot,\cdot)$. For more related results, we refer to (Hsieh et al., 2019; Solodov and Svaiter, 1999; Kotsalis et al., 2022; Ouyang and Xu, 2021) and the references therein.
@@
For nonconvex-strongly concave minimax problems, various first-order algorithms have been proposed in recent works (Bot and Böhm, 2023; Jin et al., 2020; Lin et al., 2020a, b; Lu et al., 2020), and all of them can achieve the iteration complexity of $\tilde{\mathcal{O}}(\ell\kappa_{y}^{2}\epsilon^{-2})$ in terms of stationary point of $\Phi(\cdot)=\max_{y\in\mathcal{Y}}f(\cdot,y)$ (when $x\in\mathbb{R}^{m}$), or stationary point of $f(x,y)$, where $\kappa_{y}$ is the condition number for $f(x,\cdot)$. Furthermore, Zhang et al. (2021) proposed a generic acceleration framework which can improve the iteration complexity to $\tilde{\mathcal{O}}(\ell\sqrt{\kappa_{y}}\epsilon^{-2})$. Some second-order algorithms have also been proposed in recent works (Luo et al., 2022; Chen et al., 2023; Yao and Xu, 2024) to find second-order stationary point.
@@
For general nonconvex-concave minimax problems, there are two types of algorithms, i.e., multi-loop algorithms and single-loop algorithms. For example, various multi-loop algorithms have been proposed in (Kong and Monteiro, 2021; Lin et al., 2020b; Nouiehed et al., 2019; Ostrovskii et al., 2021; Rafique et al., 2022; Thekumparampil et al., 2019). The best-known iteration complexity of multi-loop algorithms for solving nonconvex-concave minimax problems is $\tilde{\mathcal{O}}(\ell^{1.5}\epsilon^{-2.5})$, which was achieved by Lin et al. (2020b). Various single-loop algorithms have also been proposed, e.g., the gradient descent-ascent (GDA) algorithm (Lin et al., 2020a), the hybrid block successive approximation (HiBSA) algorithm (Lu et al., 2020), the unified single-loop alternating gradient projection (AGP) algorithm (Xu et al., 2023b), and the smoothed GDA algorithm (Zhang et al., 2023). Both the AGP algorithm and the smoothed GDA algorithm achieve the best-known iteration complexity, i.e., $\mathcal{O}(\ell^{4}\epsilon^{-4})$, among single-loop algorithms for solving nonconvex-concave minimax problems. For solving nonconvex-linear minimax problems, Pan et al. (2021) proposed a new alternating gradient projection algorithm, which finds an $\epsilon$-first-order Nash equilibrium point (Nouiehed et al., 2019) within $\mathcal{O}(\ell^{2}\epsilon^{-3})$ iterations.
@@
Next, we briefly review some existing works that focus on zeroth-order algorithms for solving the minimax optimization problem. For instance, Wang et al. (2020) proposed a zeroth-order gradient descent ascent (ZO-GDA) algorithm and a zeroth-order gradient descent multi-step ascent (ZO-GDMSA) algorithm for the nonconvex-strongly concave setting, and the total complexity to find an $\epsilon$-stationary point of $\Phi(\cdot)=\max_{y\in\mathcal{Y}}f(\cdot,y)$ is $\mathcal{O}(\kappa_{y}^{5}(m+n)\epsilon^{-2})$ and $\mathcal{O}(\kappa_{y}(n+m\kappa_{y}\log(\epsilon^{-1}))\epsilon^{-2})$, respectively. Xu et al. (2024) proposed a zeroth-order alternating stochastic gradient projection (ZO-AGP) algorithm to solve the nonconvex-concave minimax problem, and proved that the iteration complexity of the ZO-AGP algorithm is $\mathcal{O}(\ell^{4}\epsilon^{-4})$ with the number of function value estimates per iteration being bounded by $\mathcal{O}(n+m)$. Xu et al. (2023a) proposed a zeroth-order alternating gradient descent ascent (ZO-AGDA) algorithm and a zeroth-order variance reduced alternating gradient descent ascent (ZO-VRAGDA) algorithm for solving a class of nonconvex-nonconcave minimax problems under the deterministic and the stochastic setting, respectively. The iteration complexity of the ZO-AGDA algorithm and the ZO-VRAGDA algorithm have been proved to be $\mathcal{O}(\ell(n+m\kappa_{y}^{2})\epsilon^{-2})$ and $\mathcal{O}(\ell\kappa_{y}^{4}nm\epsilon^{-3})$, respectively. For more related results, we refer to (Liu et al., 2020; Beznosikov et al., 2020; Sadiev et al., 2021; Xu et al., 2020; Shen et al., 2023) and the references therein.